OpenStudy (anonymous):

is this a function or relation? (3,1), (3,2), (3,3), (3,-2), (3,0)

OpenStudy (anonymous):

no it's not a function

OpenStudy (amistre64):

if we define the first components to be inputs; and the second components to be output; then this is not a function

OpenStudy (amistre64):

If one input has more than one output, its not a function by definition

OpenStudy (anonymous):

Exactly because a function is defined as being one to one

OpenStudy (amistre64):

not one to one perse; unless im misunderstanding the vernacular

OpenStudy (amistre64):

many to one is fine, one to one is invertible; but one to many is not a function

OpenStudy (anonymous):

Hmm, I thought I remebered the def being that. Oh well

OpenStudy (amistre64):

x^2 is a function, but it is not one to one (-2)^2 = (2)^2 = 4

OpenStudy (anonymous):

Ahh yes you are very much so correct, makes more sense that way :)